Find the Young's modulus of a rod of diameter 30mm and length 300 mm which is subjected to a tensile load of 60 kN and the extension of the rod is equal to 0.4mm


1
Expert's answer
2021-02-02T01:14:53-0500

Solution:


As given that,

Diameter of Rod("D") = 30mm

Length of Rod("L") = 300mm

Tensile load("T") = 60KN

Extension("\delta l") Occurred in rod due to Tensile load = 0.4mm


In general, The Extension of Rod which is subjected to tensile load is given by the following formula,

"\delta l=\frac {TL} {AE}"

Where,

"A=\frac {\pi} {4} D^2" = Cross sectional area of the rod

E = Young's Modulus of Rod


So that from above Equation, Young's modulus of the rod is given by

"E = \frac {TL} {T \delta l}"

"= \frac {60 \times 10^3 \times 3 } {\frac {\pi} {4} \times0.03^2 \times 0.4 \times 10^{-3}} \\
= \frac {18 \times 10^{3}} {0.0002826 \times 10^{-3}} \\
= 63.694 \times 10^9 N/m^2 \\
=63.6 \times 10^{9}N/m^2 \\
=63.6 GN/m^2"


Young's modulus of the rod = 63.6 GN/m2